OpenStudy (anonymous):

Find an exact value. sin(17pi/12)

OpenStudy (anonymous):

I have to use pen on paper..

OpenStudy (anonymous):

I keep getting this: \[\frac{ -\sqrt{6} - 2 }{ 4 }\]

OpenStudy (anonymous):

Is that right?

OpenStudy (anonymous):

It can be.. I can't say without trying..

OpenStudy (brooke2353):

4.448 with three repeating

OpenStudy (anonymous):

That's not an option. They want an answer like the one I came up with^^

OpenStudy (brooke2353):

okay give me a minute

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

May not be that..

OpenStudy (anonymous):

\[\frac{1 + \sqrt{3}}{2 \sqrt{2}}\]

OpenStudy (brooke2353):

= [sqrt(2) - sqrt(6)] / 4 thats how it is in my answers to every math problem book

OpenStudy (anonymous):

That's right! Thanks Brooke!

OpenStudy (brooke2353):

your welcome

OpenStudy (anonymous):

I think it should be: \[\frac{2 + \sqrt{6}}{4}\]

OpenStudy (anonymous):

Leave it, if you got it right, then Good..